### v2

# X-Dreamer: Creating High-quality 3D Content by Bridging the Domain Gap Between Text-to-2D and Text-to-3D Generation

Abstract
In recent times, automatic text-to-3D content creation has made significant progress, driven by the development of pretrained 2D diffusion models. Existing text-to-3D methods typically optimize the 3D representation to ensure that the rendered image aligns well with the given text, as evaluated by the pretrained 2D diffusion model. Nevertheless, a substantial domain gap exists between 2D images and 3D assets, primarily attributed to variations in camera-related attributes and the exclusive presence of foreground objects. Consequently, employing 2D diffusion models directly for optimizing 3D representations may lead to suboptimal outcomes. To address this issue, we present X-Dreamer, a novel approach for high-quality text-to-3D content creation that effectively bridges the gap between text-to-2D and text-to-3D synthesis. The key components of X-Dreamer are two innovative designs: Camera-Guided Low-Rank Adaptation (CG-LoRA) and Attention-Mask Alignment (AMA) Loss. CG-LoRA dynamically incorporates camera information into the pretrained diffusion models by employing camera-dependent generation for trainable parameters. This integration enhances the alignment between the generated 3D assets and the camera’s perspective. AMA loss guides the attention map of the pretrained diffusion model using the binary mask of the 3D object, prioritizing the creation of the foreground object. This module ensures that the model focuses on generating accurate and detailed foreground objects. Extensive evaluations demonstrate the effectiveness of our proposed method compared to existing text-to-3D approaches. Our project webpage: https://xmu-xiaoma666.github.io/Projects/X-Dreamer.
1 Introduction
The field of text-to-3D synthesis, which seeks to generate superior 3D content predicated on input textual descriptions, has shown significant potential to impact a diverse range of applications. These applications extend beyond traditional areas such as architecture, animation, and gaming, and encompass contemporary domains like virtual and augmented reality.
Figure 1: The outputs of the text-to-2D generation model (left) and the text-to-3D generation model (right) under the same text prompt, i.e., “A statue of Leonardo DiCaprio’s head”.
In recent years, extensive research has demonstrated significant performance improvement in the text-to-2D generation task [24, 26, 25, 2] by leveraging pretrained diffusion models [30, 6, 31] on a large-scale text-image dataset [27]. Building on these advancements, DreamFusion [22] introduces an effective approach that utilizes a pretrained 2D diffusion model [26] to autonomously generate 3D assets from text, eliminating the need for a dedicated 3D asset dataset. A key innovation introduced by DreamFusion is the Score Distillation Sampling (SDS) algorithm. This algorithm aims to optimize a single 3D representation, such as NeRF [18], to ensure that rendered images from any camera perspective maintain a high likelihood with the given text, as evaluated by the pretrained 2D diffusion model. Inspired by the groundbreaking SDS algorithm, several recent works [13, 4, 16, 35, 36] have emerged, envisioning the text-to-3D generation task through the application of pretrained 2D diffusion models.
While text-to-3D generation has made significant strides through the utilization of pretrained text-to-2D diffusion models, it is crucial to recognize and address the persistent and substantial domain gap that remains between text-to-2D and text-to-3D generation. This distinction is clearly illustrated in Fig. 1. To begin with, the text-to-2D model produces camera-independent generation results, focusing on generating high-quality images from specific angles while disregarding other angles. In contrast, 3D content creation is intricately tied to camera parameters such as position, shooting angle, and field of view. As a result, a text-to-3D model must generate high-quality results across all possible camera parameters. This fundamental difference emphasizes the necessity for innovative approaches that enable the pretrained diffusion model to consider camera parameters. Furthermore, a text-to-2D generation model must simultaneously generate both foreground and background elements while maintaining the overall coherence of the image. Conversely, a text-to-3D generation model only needs to concentrate on creating the foreground object. This distinction allows text-to-3D models to allocate more resources and attention to precisely represent and generate the foreground object. Consequently, the domain gap between text-to-2D and text-to-3D generation poses a significant performance obstacle when directly employing pretrained 2D diffusion models for 3D asset creation.
In this study, we present a pioneering framework, X-Dreamer, designed to address the domain gap between text-to-2D and text-to-3D generation, thereby facilitating the creation of high-quality text-to-3D content. Our framework incorporates two innovative designs that are specifically tailored to address the aforementioned challenges. Firstly, existing approaches [13, 4, 16, 35] commonly employ 2D pretrained diffusion models [26, 25] for text-to-3D generation, which lack inherent linkage to camera parameters. To address this limitation and ensure that our text-to-3D model produces results that are directly influenced by camera parameters, we introduce Camera-Guided Low-Rank Adaptation (CG-LoRA) to fine-tune the pretrained 2D diffusion model. Notably, the parameters of CG-LoRA are dynamically generated based on the camera information during each iteration, establishing a robust relationship between the text-to-3D model and camera parameters. Furthermore, pretrained text-to-2D diffusion models allocate attention to both foreground and background generation, whereas the creation of 3D assets necessitates a stronger focus on accurately generating foreground objects. To address this requirement, we introduce Attention-Mask Alignment (AMA) Loss, which leverages the rendered binary mask of the 3D object to guide the attention map of the pretrained 2D stable diffusion model [25]. By incorporating this module, X-Dreamer prioritizes the generation of foreground objects, resulting in a significant enhancement of the overall quality of the generated 3D content.
We present a compelling demonstration of the effectiveness of X-Dreamer in synthesizing high-quality 3D assets based on textual cues. By incorporating CG-LoRA and AMA loss to address the domain gap between text-to-2D and text-to-3D generation, our proposed framework exhibits substantial advancements over prior methods in text-to-3D generation. In summary, our study contributes to the field in three key aspects:
  • We propose a novel method, X-Dreamer, for high-quality text-to-3D content creation, effectively bridging the domain gap between text-to-2D and text-to-3D generation.
  • To enhance the alignment between the generated results and the camera perspective, we propose CG-LoRA, which leverages camera information to dynamically generate specific parameters for 2D diffusion models.
  • To prioritize the creation of foreground objects in the text-to-3D model, we introduce AMA loss, which utilizes binary masks of the foreground 3D object to guide the attention maps of the 2D diffusion model.
2 Related Work
Figure 2: Overview of the proposed X-Dreamer, which consists of geometry learning and appearance learning.
2.1 Text-to-3D Content Creation
In recent years, there has been a significant surge in interest surrounding the evolution of text-to-3D generation [17, 22, 12]. This growing field has been propelled, in part, by advancements in pretrained vision-and-language models, such as CLIP [23], as well as diffusion models like Stable Diffusion [25] and Imagen [26]. Contemporary text-to-3D models can generally be classified into two distinct categories: the CLIP-based text-to-3D approach and the diffusion-based text-to-3D approach. The CLIP-based text-to-3D approach [17, 14, 11, 9, 19, 37] employs CLIP encoders [23] to project textual descriptions and rendered images derived from the 3D object into a modal-shared feature space. Subsequently, CLIP loss is harnessed to align features from both modalities, optimizing the 3D representation to conform to the textual description. Various scholars have made significant contributions to this field. For instance, Michel et al. [17] are pioneers in proposing the use of CLIP loss to harmonize the text prompt with the rendered images of the 3D object, thereby enhancing text-to-3D generation. Ma et al. [14] introduce dynamic textual guidance during 3D object synthesis to improve convergence speed and generation performance. However, these approaches have inherent limitations, as they tend to generate 3D representations with a scarcity of geometry and appearance detail. To overcome this shortcoming, the diffusion-based text-to-3D approach [22, 13, 4, 32, 10, 8] leverages pretrained text-to-2D diffusion models [25, 26] to guide the optimization of 3D representations. Central to these models is the application of SDS loss [22] to align the rendered images stemming from a variety of camera perspectives with the textual description. Specifically, given the target text prompt, Lin et al. [13] leverage a coarse-to-fine pipeline to generate high-resolution 3D content. Chen et al. [4] decouple geometry modeling and appearance modeling to generate realistic 3D assets. For specific purposes, some researchers [28, 36] integrate trainable LoRA [7] branches into pretrained diffusion models. For instance, Seo et al. [28] put forth 3DFuse, a model that harnesses the power of LoRA to comprehend object semantics. Wang et al. [36] introduce ProlificDreamer, where the role of LoRA is to evaluate the score of the variational distribution for 3D parameters. However, the LoRA parameter begins its journey from random initialization and maintains its independence from the camera and text. To address these limitations, we present two innovative modules: CG-LoRA and AMA loss. These modules are designed to enhance the model’s ability to consider important camera parameters and prioritize the generation of foreground objects throughout the text-to-3D creation process.
2.2 Low-Rank Adaptation (LoRA)
Low-Rank Adaptation (LoRA) is a technique used to reduce memory requirements when fine-tuning a large model. It involves injecting only a small set of trainable parameters into the pretrained model, while keeping the original parameters fixed. During the optimization process, gradients are passed through the fixed pretrained model weights to the LoRA adapter, which is then updated to optimize the loss function. LoRA has been applied in various fields, including natural language processing [7, 3], image synthesis [38] and 3D generation [28, 36]. To achieve low-rank adaptation, a linear projection with a pretrained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{in}\times d_{out}}$ is augmented with an additional low-rank factorized projection. This augmentation is represented as $\mathbf{W}_{0}+\Delta\mathbf{W}=\mathbf{W}_{0}+\mathbf{A}\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{d_{in}\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times d_{out}}$, and $r\ll\min(d_{in},d_{out})$. During training, $\mathbf{W}_{0}$ remains fixed, while $\mathbf{A}$ and $\mathbf{B}$ are trainable. The modified forward pass, given the original forward pass $\mathbf{Y}=\mathbf{X}\mathbf{W}_{0}$, can be formulated as follows:
$\mathbf{Y}=\mathbf{X}\mathbf{W}_{0}+\mathbf{X}\mathbf{A}\mathbf{B}.$ (1)
In this paper, we introduce CG-LoRA, which involves the dynamic generation of trainable parameters for $\mathbf{A}$ based on camera information. This technique allows for integrating perspective information, including camera parameters and direction-aware descriptions, into the pretrained text-to-2D diffusion model. As a result, our method significantly enhances text-to-3D generation capabilities.
3 Approach
Figure 3: Illustration of Camera-Guided Low-Rank Adaptation.
3.1 Architecture
In this section, we present a comprehensive introduction to the proposed X-Dreamer, which consists of two main stages: geometry learning and appearance learning. For geometry learning, we employ DMTet [29] as the 3D representation. DMTet is an MLP parameterized with $\Phi_{dmt}$ and is initialized with a 3D ellipsoid using the mean squared error (MSE) loss $\mathcal{L}_{MSE}$. Subsequently, we optimize DMTet and CG-LoRA using the SDS loss [22] $\mathcal{L}_{SDS}$ and the proposed AMA loss $\mathcal{L}_{AMA}$ to ensure the alignment between the 3D representation and the input text prompt. For appearance learning, we leverage bidirectional reflectance distribution function (BRDF) modeling [33] following the previous approach [4]. Specifically, we utilize an MLP with trainable parameters $\Phi_{mat}$ to predict surface materials. Similar to the geometry learning stage, we optimize $\Phi_{mat}$ and CG-LoRA using the SDS loss $\mathcal{L}_{SDS}$ and the AMA loss $\mathcal{L}_{AMA}$ to achieve alignment between the 3D representation and the text prompt. Fig. 2 provides a detailed depiction of our proposed X-Dreamer.
Geometry Learning. For geometry learning, an MLP network $\Phi_{dmt}$ is utilized to parameterize DMTet as a 3D representation. To enhance the stability of geometry modeling, we employ a 3D ellipsoid as the initial configuration for DMTet $\Phi_{dmt}$. For each vertex $v_{i}\in V_{T}$ belonging to the tetrahedral grid $T$, we train $\Phi_{dmt}$ to predict two important values: the SDF value $s(v_{i})$ and the deformation offset $\delta(v_{i})$. To initialize $\Phi_{dmt}$ with the 3D ellipsoid, we sample a set of $N$ points $\{p_{i}\in\mathbb{R}^{3}\}|_{i=1}^{N}$ approximately distributed on the surface of an ellipsoid and compute the corresponding SDF values $\{{SDF}(p_{i})\}|_{i=1}^{N}$. Subsequently, we optimize $\Phi_{dmt}$ using MSE loss. This optimization process ensures that $\Phi_{dmt}$ effectively initializes DMTet to resemble the 3D ellipsoid. The formulation of the MSE loss is given by:
$\mathcal{L}_{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(s(p_{i};\Phi_{dmt})-SDF(p_{i})\right)^{2}.$ (2)
After initializing the geometry, our objective is to align the geometry of DMTet with the input text prompt. Specifically, we generate the normal map $\boldsymbol{n}$ and the object mask $\boldsymbol{m}$ from the initialized DMTet $\Phi_{dmt}$ by employing a differentiable rendering technique [33], given a randomly sampled camera pose $\boldsymbol{c}$. Subsequently, we input the normal map $\boldsymbol{n}$ into the frozen stable diffusion (SD) with a trainable CG-LoRA and update $\Phi_{dmt}$ using the SDS loss, which is defined as follows:
$\small\nabla_{\Phi_{dmt}}\mathcal{L}_{\mathrm{SDS}}=\mathbb{E}_{t,\boldsymbol{\epsilon}}\left[w(t)\left(\hat{\boldsymbol{\epsilon}}_{\Theta}(\boldsymbol{n}_{t};\boldsymbol{y},t)-\boldsymbol{\epsilon}\right)\frac{\partial\boldsymbol{n}}{\partial{\Phi_{dmt}}}\right],$ (3)
where $\Theta$ represents the parameter of SD, $\hat{\boldsymbol{\epsilon}}_{\Theta}(\boldsymbol{n}_{t};\boldsymbol{y},t)$ denotes the predicted noise of SD given the noise level $t$ and text embedding $\boldsymbol{y}$. Additionally, $\boldsymbol{n}_{t}=\alpha_{t}\boldsymbol{n}+\sigma_{t}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{I})$ represents noise sampled from a normal distribution. The implementation of $w(t)$, $\alpha_{t}$, and $\sigma_{t}$ is based on the DreamFusion [22].
Furthermore, to focus SD on generating foreground objects, we introduce an additional AMA loss to align the object mask $\boldsymbol{m}$ with the attention map of SD, given by:
$\mathcal{L}_{AMA}=\frac{1}{L}\sum_{i=1}^{L}|\boldsymbol{a}_{i}-\eta(\boldsymbol{m})|,$ (4)
where $L$ denotes the number of attention layers, and $\boldsymbol{a}_{i}$ is the attention map of $i$-th attention layer. The function $\eta(\cdot)$ is employed to resize the rendered mask, ensuring its dimensions align with those of the attention maps.
Appearance Learning. After obtaining the geometry of the 3D object, our objective is to compute its appearance using the Physically-Based Rendering (PBR) material model [15]. The material model comprises the diffuse term $\boldsymbol{k}_{d}\in\mathbb{R}^{3}$, the roughness and metallic term $\boldsymbol{k}_{rm}\in\mathbb{R}^{2}$, and the normal variation term $\boldsymbol{k}_{n}\in\mathbb{R}^{3}$. For any point $p\in\mathbb{R}^{3}$ on the surface of the geometry, we utilize an MLP parameterized by $\Phi_{mat}$ to obtain the three material terms, which can be expressed as follows:
$(\boldsymbol{k}_{d},\boldsymbol{k}_{n},\boldsymbol{k}_{rm})=\text{MLP}\left(\mathcal{P}(p);\Phi_{mat}\right),$ (5)
where $\mathcal{P}(\cdot)$ represents the positional encoding using a hash-grid technique [20]. Subsequently, each pixel of the rendered image can be computed as follows:
$V(p,\omega)=\int_{\Omega}L_{i}(p,\omega_{i})f(p,\omega_{i},\omega)(\omega_{i}\cdot n_{p})\mathrm{d}\omega_{i},$ (6)
where $V(p,\omega)$ denotes the rendered pixel value from the direction $\omega$ for the surface point $p$. $\Omega$ denotes a hemisphere defined by the set of incident directions $\omega_{i}$ satisfying the condition $\omega_{i}\cdot n_{p}\geq 0$, where $\omega_{i}$ denotes the incident direction, and $n_{p}$ represents the surface normal at point $p$. $L_{i}(\cdot)$ corresponds to the incident light from an off-the-shelf environment map, and $f(\cdot)$ is the Bidirectional Reflectance Distribution Function (BRDF) related to the material properties (i.e., $\boldsymbol{k}_{d},\boldsymbol{k}_{n},\boldsymbol{k}_{rm}$). By aggregating all rendered pixel colors, we obtain a rendered image $\boldsymbol{x}=\{V(p,\omega)\}$. Similar to the geometry modeling stage, we feed the rendered image $\boldsymbol{x}$ into SD. The optimization objective remains the same as Eq. 3 and Eq. 4, where the rendered normal map $\boldsymbol{n}$ and the parameters of DMTet $\Phi_{det}$ are replaced with the rendered image $\boldsymbol{x}$ and the parameters of the material encoder $\Phi_{mat}$, respectively.
3.2 Camera-Guided Low-Rank Adaptation
The domain gap between text-to-2D and text-to-3D generation presents a significant challenge, as discussed in Sec. 1. It has been observed that directly utilizing pretrained SD for text-to-3D generation can result in certain issues, such as the Janus Problem [12, 1]. To address these issues, we propose Camera-Guided Low-Rank Adaptation (CG-LoRA) as a solution to bridge the domain gap. As depicted in Fig. 20, we leverage camera parameters and direction-aware text to guide the generation of parameters in CG-LoRA, enabling X-Dreamer to effectively incorporate camera perspective and direction information.
Specifically, given a text prompt $T$ and camera parameters $C=\{x,y,z,\phi_{yaw},\phi_{pit},\theta_{fov}\}$ 111 The variables $x,y,z,\phi_{yaw},\phi_{pit},\theta_{fov}$ represent the x, y, z coordinates, yaw angle, pitch angle of the camera, and field of view, respectively. The roll angle $\phi_{roll}$ is intentionally set to 0 to ensure the stability of the object in the rendered image., we initially project these inputs into a feature space using the pretrained textual CLIP encoder $\mathcal{E}_{txt}(\cdot)$ and a trainable MLP $\mathcal{E}_{pos}(\cdot)$:
$\boldsymbol{t}$ $=\mathcal{E}_{txt}(T),$ (7)
$\boldsymbol{c}$ $=\mathcal{E}_{pos}(C),$ (8)
where $\boldsymbol{t}\in\mathbb{R}^{d_{txt}}$ and $\boldsymbol{c}\in\mathbb{R}^{d_{cam}}$ are textual features and camera features. Subsequently, we employ two low-rank matrices to project $\boldsymbol{t}$ and $\boldsymbol{c}$ into trainable dimensionality-reduction matrices within CG-LoRA:
$\boldsymbol{A}_{txt}$ $=\text{Reshape}(\boldsymbol{t}\boldsymbol{W}_{txt}),$ (9)
$\boldsymbol{A}_{cam}$ $=\text{Reshape}(\boldsymbol{c}\boldsymbol{W}_{cam}),$ (10)
[…]
$\boldsymbol{y}=\boldsymbol{x}\boldsymbol{W}+[\boldsymbol{x}\boldsymbol{A}_{txt};\boldsymbol{x}\boldsymbol{A}_{cam}]\boldsymbol{B},$ (11)
where $\boldsymbol{W}\in\mathbb{R}^{d\times d}$ represents the frozen parameters of the pretrained SD model, and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is the concatenation operation alone the channel dimension. In our implementation, we integrate CG-LoRA into the linear embedding layers of the attention modules in SD to effectively capture direction and camera information.
Figure 4: Text-to-3D generation results from an ellipsoid.
Figure 5: Text-to-3D generation results from coarse-grained guided meshes.
Figure 6: Comparison with State-of-the-Art (SOTA) methods. Our method yields results that exhibit enhanced fidelity and more details.
3.3 Attention-Mask Alignment Loss
Although SD is pretrained to generate 2D images that encompass both foreground and background elements, the task of text-to-3D generation demands a stronger focus on generating foreground objects. To address this specific requirement, we introduce Attention-Mask Alignment (AMA) Loss, which aims to align the attention map of SD with the rendered mask image of the 3D object. Specifically, for each attention layer in the pretrained SD, we compute the attention map between the query image feature $\boldsymbol{Q}\in\mathbb{R}^{H\times h\times w\times\frac{d}{H}}$ and the key CLS token feature $\boldsymbol{K}\in\mathbb{R}^{H\times\frac{d}{H}}$. The calculation is formulated as follows:
$\bar{\boldsymbol{a}}=\mathrm{Softmax}(\frac{\boldsymbol{Q}\boldsymbol{K}^{\top}}{\sqrt{d}}),$ (12)
where $H$ denotes the number of attention heads in SD, and $\bar{\boldsymbol{a}}\in\mathbb{R}^{H\times h\times w}$ represents the attention map. Subsequently, we proceed to compute the overall attention map $\hat{\boldsymbol{a}}\in\mathbb{R}^{h\times w}$ by averaging the attention values of $\bar{\boldsymbol{a}}$ across all attention heads. Since the attention map values are normalized using the softmax function, the activation values in the attention map may become very small when the image feature resolution is high. However, considering that each element in the rendered mask has a binary value of either 0 or 1, directly aligning the attention map with the rendered mask is not optimal. To address this, we propose a normalization technique that maps the values in the attention map from 0 to 1. This normalization process is formulated as follows:
$\boldsymbol{a}=\frac{\hat{\boldsymbol{a}}-\min(\hat{\boldsymbol{a}})}{\max(\hat{\boldsymbol{a}})-\min(\hat{\boldsymbol{a}})+\nu},$ (13)
where $\nu$ represents a small constant value (e.g., $-61e\textrm{-}$ 6) that prevents division by zero in the denominator. Finally, we align the attention maps of all attention layers with the rendered mask of the 3D object using the AMA loss. The formulation of this alignment is presented in Eq. 4.
4 Experiments
4.1 Implementation Details.
We conduct the experiments using four Nvidia RTX 3090 GPUs and the PyTorch library [21]. To calculate the SDS loss, we utilize the Stable Diffusion implemented by HuggingFace Diffusers [34]. For the DMTet $\Phi_{dmt}$ and material encoder $\Phi_{mat}$, we implement them as a two-layer MLP and a single-layer MLP, respectively, with a hidden dimension of 32. The values of $d_{cam}$, $d_{txt}$, $r$, $r^{\prime}$, the batch size, the SDS loss weight, the AMA loss weight, and the aspect ratio of the perspective projection plane are set to 1024, 1024, 4, 4, 4, 1, 0.1, and 1 respectively. We optimize X-Dreamer for 2000 iterations for geometry learning and 1000 iterations for appearance learning. For each iteration, $\phi_{pit}$, $\phi_{yaw}$, and $\theta_{fov}$ are randomly sampled from $(-15^{\circ},45^{\circ})$, $(-180^{\circ},180^{\circ})$, and $(25^{\circ},45^{\circ})$, respectively.
4.2 Results of X-Dreamer
Text-to-3D generation from an ellipsoid. We present representative results of X-Dreamer for text-to-3D generation, utilizing an ellipsoid as the initial geometry, as shown in Fig. 4. The results demonstrate the ability of X-Dreamer to generate high-quality and photo-realistic outputs that accurately correspond to the input text prompts.
Text-to-3D generation from coarse-grained meshes. While there is a wide availability of coarse-grained meshes for download from the internet, directly utilizing these meshes for 3D content creation often results in poor performance due to the lack of geometric details. However, when compared to a 3D ellipsoid, these meshes may provide better 3D shape prior information for X-Dreamer. Hence, instead of using ellipsoids, we can initialize DMTet with coarse-grained guided meshes as well. As shown in Fig. 5, X-Dreamer can generate 3D assets with precise geometric details based on the given text, even when the provided coarse-grained mesh lacks details. For instance, in the last column of Fig. 5, X-Dreamer accurately transforms the geometry from a cow to a corgi based on the text prompt “A corgi, highly detailed.” Therefore, X-Dreamer is also an exceptionally powerful tool for editing coarse-grained mesh geometry using textual inputs.
Figure 7: Ablation studies of the proposed X-Dreamer.
Figure 8: Visualization of Attention Map, Rendered Mask, and Rendered Image with and without AMA Loss. For clarity, we only visualize the attention map of the first attention layer in SD.
Qualitative Comparison. To assess the effectiveness of X-Dreamer, we compare it with four SOTA methods: DreamFusion [22], Magic3D [13], Fantasia3D [4], and ProlificDreamer [36], as depicted in Fig. 6. When compared to the SDS-based methods [22, 13, 36], X-Dreamer outperforms them in generating superior-quality and realistic 3D assets. In addition, when compared to the VSD-based method [36], X-Dreamer produces 3D content with comparable or even better visual effects, while requiring significantly less optimization time. Specifically, the geometry and appearance learning process of X-Dreamer requires only approximately 27 minutes, whereas ProlificDreamer exceeds 8 hours.
4.3 Ablation Study
Ablation on the proposed modules. To gain insights into the abilities of CG-LoRA and AMA loss, we perform ablation studies wherein each module is incorporated individually to assess its impact. As depicted in Fig. 7, the ablation results demonstrate a notable decline in the geometry and appearance quality of the generated 3D objects when CG-LoRA is excluded from X-Dreamer. For instance, as shown in the second row of Fig. 7, the generated Batman lacks an ear on the top of its head in the absence of CG-LoRA. This observation highlights the crucial role of CG-LoRA in injecting camera-relevant information into the model, thereby enhancing the 3D consistency. Furthermore, the omission of AMA loss from X-Dreamer also has a deleterious effect on the geometry and appearance fidelity of the generated 3D assets. Specifically, as illustrated in the first row of Fig. 7, X-Dreamer successfully generates a photorealistic texture for the rocket, whereas the texture quality noticeably deteriorates in the absence of AMA loss. This disparity can be attributed to AMA loss, which directs the focus of the model towards foreground object generation, ensuring the realistic representation of both geometry and appearance of foreground objects. These ablation studies provide valuable insights into the individual contributions of CG-LoRA and AMA loss in enhancing the geometry, appearance, and overall quality of the generated 3D objects.
Attention map comparisons w/ and w/o AMA loss. AMA loss is introduced with the aim of guiding attention during the denoising process towards the foreground object. This objective is achieved by aligning the attention map of SD with the rendered mask of the 3D object. To evaluate the effectiveness of AMA loss in accomplishing this goal, we visualize the attention maps of SD with and without AMA loss at both the geometry learning and appearance learning stages. As depicted in Fig. 8, it can be observed that incorporating AMA loss not only results in improved geometry and appearance of the generated 3D asset, but also concentrates the attention of SD specifically on the foreground object area. The visualizations confirm the efficacy of AMA loss in directing the attention of SD, resulting in improved quality and foreground object focus during geometry and appearance learning stages.
5 Conclusion
This study introduces a groundbreaking framework called X-Dreamer, which is designed to enhance text-to-3D synthesis by addressing the domain gap between text-to-2D and text-to-3D generation. To achieve this, we first propose CG-LoRA, a module that incorporates 3D-associated information, including direction-aware text and camera parameters, into the pretrained Stable Diffusion (SD) model. By doing so, we enable the effective capture of information relevant to the 3D domain. Furthermore, we design AMA loss to align the attention map generated by SD with the rendered mask of the 3D object. The primary objective of AMA loss is to guide the focus of the text-to-3D model towards the generation of foreground objects. Through extensive experiments, we have thoroughly evaluated the efficacy of our proposed method, which has consistently demonstrated its ability to synthesize high-quality and photorealistic 3D content from given text prompts.